OpenStudy (anonymous):

1/5 questions to check, ty! Jorge bought a new computer for $399 and extra computer memory for $42. What percent of the total amount that Jorge spent was the cost of the computer? Round to the nearest tenth of a percent, if necessary. A. 9.5% B. 10.5% C. 90.5% D. 92.6% >>>>>90.5%<<<<<

OpenStudy (aveline):

Correct

OpenStudy (anonymous):

Okay, thanks! Can you check 4 more please? @Aveline

OpenStudy (aveline):

Yup

OpenStudy (anonymous):

Alright, should I make another question thing?

OpenStudy (anonymous):

Joseph bought some woodworking equipment. He paid $62 for a saw, $9 for a file set, and $4 for sandpaper. What percent of the total cost of the equipment that Joseph bought was the price of the file set? Round to the nearest percent, if necessary. 833%?

OpenStudy (aveline):

9/75 is definitely not 833%

OpenStudy (anonymous):

Yeah true, I think I switched them around (I put 75/9)

OpenStudy (anonymous):

Is it 6.25?

OpenStudy (aveline):

Just plug in 9/75 into your calculator/google. Then multiply by 100 It should be 12%

OpenStudy (anonymous):

oh, alright next one then. Thanks so far! Kwame purchased a washer for $312.20. The price of the washer included sales tax of 11.5%. Which equation can be used to find the price, p, of the washer before tax? A. 11.5p = 312.20 B. p + 0.115p = 312.20 C. 0.115p = 312.20 D. p + 11.5p = 312.20 B?

OpenStudy (aveline):

Yes

OpenStudy (anonymous):

Okay, 2 more!! Billie’s monthly cell phone bill includes the cost of cell phone service, 15% tax on the cost of the cell phone service, and a $9.25 data fee. Billie’s monthly cell phone bill totals $61. What is the cost of the cell phone service before tax is added? 45.00?

OpenStudy (aveline):

Yupperdoodles

OpenStudy (anonymous):

Yay! 1 more :D At a car wash, Lily paid $11.75 for the wash, 8% sales tax on the price of the wash, and $2 for the tip. Which statements are true? Choose exactly two answers that are correct. A. The equation that represents the total price, w, of the car wash is w + 0.08w + 2 = 11.75. B. The equation that represents the total price, w, of the car wash is 11.75 + 0.08(11.75) + 2 = w. C. The total cost of the car wash is $9.03. D. The total cost of the car wash is $14.69. ~~~C and A?~~~

OpenStudy (aveline):

A is correct, but not C. If Lily paid $11.75 for the car wash, plus two additional fees, how can it cost $9.03? ...assuming that Lily is not an extreme couponer ;)

OpenStudy (anonymous):

Oh, but I added all of it and it came out as 13.83 How do I get $14.69?

OpenStudy (aveline):

11.75*1.08=12.69 12.69+2.00=$14.69

OpenStudy (anonymous):

Oh, I see. One more question, how does 0.08 change into 1.08?

OpenStudy (aveline):

The car wash has a sales tax of 8%. Before tax, the price is 11.75. If we want to find how much the tax costs, we can do a variety of things: We can multiply 11.75*0.08=0.94 Then, we can add this to 11.75 0.94+11.75=12.69 We can also just multiply 11.75 by 1.08, or 108%, like I did previously

OpenStudy (kittiwitti1):

☻ What's going on right meow?

OpenStudy (aveline):

@kittiwitti1 Just checking some answers

OpenStudy (anonymous):

Thanks so much Aveline you're awesome! And hi Plagg!

OpenStudy (kittiwitti1):

*inside joke with @LoveIt sorry lmao*

OpenStudy (aveline):

:P

OpenStudy (aveline):

No problem Lovelt!

OpenStudy (kittiwitti1):

Hiyo Tikki, how are you doing this purrfect night~ ♥

OpenStudy (anonymous):

Wonderful, you? :D

OpenStudy (kittiwitti1):

(Terrible. Naggy mother and all > <) My day's been amewsing. Silly people n all ☻

OpenStudy (anonymous):

Awesome! (sorry about your mother lol) Well, I got to go bye Plagg~! ^_^

OpenStudy (kittiwitti1):

Bye Tikki, have a magnifique day ♥

OpenStudy (anonymous):

Bye!